### v1

# OpenACE: An Open Benchmark for Evaluating Audio Coding Performance

Abstract
Audio and speech coding lack unified evaluation and open-source testing. Many candidate systems were evaluated on proprietary, non-reproducible, or small data, and machine learning-based codecs are often tested on datasets with similar distributions as trained on, which is unfairly compared to digital signal processing-based codecs that usually work well with unseen data. This paper presents a full-band audio and speech coding quality benchmark with more variable content types, including traditional open test vectors. An example use case of audio coding quality assessment is presented with open-source Opus, 3GPP’s EVS, and recent ETSI’s LC3 with LC3+ used in Bluetooth LE Audio profiles. Besides, quality variations of emotional speech encoding at 16 kbps are shown. The proposed open-source benchmark contributes to audio and speech coding democratization and is available at https://github.com/JozefColdenhoff/OpenACE.
Index Terms:
audio coding, benchmarks, deep learning, speech processing
I Introduction
Audio and speech coding is a key component of communication and streaming systems used ubiquitously worldwide. Model-based compression methods developed decades ago and continuously improved, extensively updated, and tested by telecom companies and certification institutions have faced a paradigm shift in the last decade due to data-driven approaches fueled by machine learning (ML). Traditional DSP (Digital Signal Processing) and neural codecs are now combined into hybrid systems [1] and are often assessed together.
Audio and speech coding is democratized by recent competitive neural codec development, released by various small research and startup teams, using available open-source or popular licensed data. There are several challenges in properly evaluating new neural audio codecs: (i) valid comparison to standardized codecs that are extensively tested but usually with proprietary test sequences, (ii) valid comparison between the new codecs themselves coming from different teams, and (iii) comparisons focusing either on achieved audio quality, delay or computational efficiency. Table I outlines the databases used to evaluate recent neural codecs. First, almost every codec was evaluated using different data. Second, some datasets have restricted access; third, most contain only English language sequences. Often, the neural codecs are not tested cross-datasets, which is not fair compared to the DSP-based codecs, which are not optimized for particular data.
TABLE I: Neural codecs and their evaluation speech data. SoundStream uses also Freesound [2] noises and music from MagnaTagATune [3]. DAPS evaluates also on music from MUSDB dataset [4] N/O: Non-Open.
Codec Dataset Access Language
DAC [5] AudioSet [6] Open English
DAPS [7] Open English
SoundStream [8] VCTK [9] Open English
LibriTTS [10] Open English
Lyra [11] VCTK Open English
Zhen et al. [12] TIMIT [13] N/O English
WaveNet [14] WSJ [] N/O English
LPCNet [15] NTT [16] N/O Various
Gârbacea et al. [17] LS [18] Open English
Kankanahalli [19] TIMIT N/O English
Early attempts (2010–2016) of ML-based speech coding focused on binary quantization of speech representations, e.g. [20, 21, 22], have been followed by recent (2017–) end-to-end neural systems that jointly learn encoder, quantizer and decoder [19, 23, 24, 11, 8]. These new neural audio and speech codecs are often compared to traditional, DSP-based approaches, like open-source Opus [25], 3GPP’s EVS [26] and ETSI’s LC3/LC3+ codecs. It is worth mentioning that legacy or neural codecs usually code single audio channels. Stereo or multi-channel streams are assembled by aggregating single audio channel instances. Multi-channel codecs, like [27], also appear and this presented proposal is also valid for them.
This paper proposes the Open Audio Coding Evaluation (OpenACE) benchmark, an open-source audio and speech coding evaluation that aims to democratize novel neural or hybrid audio coding and contribute to better evaluation of modern codecs. We design and describe a joint audio and speech coding benchmark that unifies various content types, such as speech and music. We release the benchmark with an open-source evaluation code to further facilitate its use.
We provide two examples of codec comparisons. The first one compares Opus and EVS with two recent LC3 and LC3+ codecs used in Bluetooth Low Energy (BLE) Audio profiles, at varying bitrates using objective intrusive metrics. LC3 was designed for hearing aids and very low power devices. Although BLE Audio gains popularity in consumer electronics, LC3 assessment has been done yet only on small proprietary data. The second example focuses on the emotional speech encoding at 16 kbps, where we show high encoded speech quality dependency on particular emotions.
TABLE II: The OpenACE description and statistics. AE: North American English, AS: American Spanish, BE: British English, GR:German, MA:Mandarin, FR:French, DU:Dutch, FI:Finnish, IT:Italian, JP:Japanese, PL:Polish, (B/C/D):ITU-T P.501 Clause, m: male, f: female, a:artificial. Overall, the OpenACE test contains 5.9 hours from 77 speakers with 10 languages.
Dataset Languages Samp. freq. Bit Speakers Samples Av. duration
[kHz] depth [s]
IEEE 269-2010 [28] AE 44.1/48 16 4m, 4f, a 6 7.4
ETSI TS 103-281 [29] AE,GR,MA 48 16 10m, 10f 3 81
ITU-T P.501 [30] MA,DU,BE,AE,FR,GR,IT,JP,PL,AS 8/16/48 16 20m, 20f B:48 B:9
C:124 C:8
D:30 D:6
VCTK test [9] English with various accents 48 16 3m 831 2.5
EARS test  [31] English with various accents 48 32 3m, 3f 886 15
EBU SQAM [32] - (Audio) 44.1 16 - 70 50
ODAQ [33] - (Audio) 44.1 24 - 47 11
II Benchmark proposal
II-A Design Principles
Sampling frequency and bit depth: Modern codecs can scale from narrowband (8kHz) speech to high-quality full-band (48kHz) stereo audio, with bits per audio sample ranging from 16 to 32 bits. Internal calculations in LC3 is 16 bits, and 24 bits in LC3+. Most open speech data is sampled at 16kHz, traditionally enough for automatic speech recognition. However, the proposed coding benchmark should thus contain full-band audio and speech data with higher bit depths that allow downsampling to the target evaluation setting.
Legacy and availability: 3GPP, IEFT or ETSI codecs are extensively tested but with restricted access data. Even the open-source Opus project provides the testing vectors [^1] https://opus-codec.org/testvectors to verify the codec implementation but does not provide the evaluation data. Any benchmark proposal should opt for open-source data to democratize the evaluation and reproducibility of the candidate systems. There are some available test signals from the standardization organizations that should be included in any speech coding testing. Table II lists the three such datasets from the top.
Data variability: As mentioned in the introduction, speech and audio coding are used ubiquitously worldwide and for many media types. Candidate codecs need to demonstrate robustness to i) intrinsic signal variabilities that include speakers: language, age, media type: typical and atypical speech, audio events and music, and ii) extrinsic signal variabilities including environment: noisy and reverberant, and transmission errors caused, for example, by the channel and RF interference.
Speech quality assessment: The ITU-T standards Perceptual Evaluation of Speech Quality (PESQ), Perceptual Objective Listening Quality Analysis (POLQA), and ETSI’s 3-fold Quality Evaluation of Speech in Telecommunications (3QUEST) belong to the most popular intrusive speech quality methods. Clean references are required to compute the objective scores. To perform subjective/listening test correctly, the speech files need to be calibrated to an active speech level of 79 dB SPL.
II-B Benchmark data
In this work, we propose a joint audio-speech content types quality benchmark, mixing traditional open assessment datasets with recent large, higher quality, and more variable content datasets:
  • The IEEE 269-2010 [28] standardizes methods for making laboratory measurements of electroacoustic characteristics of analog and digital telephones, handsets, and headsets. The methods may also apply to various other communications equipment, including cordless, wireless, and mobile communications devices. It includes a small available full-band dataset of 8 speakers.
  • The ETSI TS 103-281 [29] provides the test audio vectors from 20 speakers in 3 languages, designed to be used to objectively evaluate the performance of super-wideband and full band mobile terminals for speech communication in the presence of background noise.
  • The ITU-T P.501 [30] recommendation gives a wide variety of test signals, starting with low-complexity test signals up to test signals with a high degree of complexity incorporating many typical parameters of speech. It contains multi-lingual test samples from ten different languages and sampling frequency variation from 8 to 48 KHz.
  • The VCTK test samples [9] are widely used in neural or hybrid speech coding proposals, as in 2017, not many full-band high-quality data were opened for researchers. Initially designed for voice cloning experiments, it includes many utterances from fewer speakers.
  • The EARS, anechoic fullband speech dataset [31], was originally benchmarked for speech enhancement and dereverberation. It is also excellent for speech coding benchmarking, as it includes a large range of different speaking styles, including emotional speech, reading styles, non-verbal sounds, and conversational freeform speech. We included all six test speakers.
  • Two audio datasets: EBU-SQAM [32] and ODAQ [33]. The latter one includes the audio material from newly released productions by Netflix and Fraunhofer IIS at 24-bit sample resolution.
TABLE III: The VISQOL and POLQA scores computed for reference and encoded audio on the full-band subset of the benchmark data. The input files were truncated to the first 10 seconds. 95% confidence intervals are computed using 1000 bootstrap samples.
16 kbps 32 kbps 64 kbps
VISQOL POLQA VISQOL POLQA VISQOL POLQA (subset)
LC3 3.51 [3.49, 3.52] 2.98 [2.95, 3.01] 4.00 [3.99, 4.01] 4.44 [4.42, 4.46] 4.39 [4.38, 4.40] 4.64 [4.63, 4.65]
LC3+ 3.11 [3.08, 3.13] 3.23 [3.20, 3.26] 4.00 [3.99, 4.01] 4.45 [4.43, 4.47] 4.39 [4.38, 4.40] 4.64 [4.63, 4.65]
EVS 3.92 [3.90, 3.93] 4.34 [4.32, 4.35] 3.82 [3.81, 3.84] 4.37 [4.36, 4.39] 4.19 [4.17, 4.20] 4.53 [4.52, 4.55]
Opus 3.98 [3.97, 3.99] 4.03 [4.01, 4.05] 4.05 [4.04, 4.06] 4.56 [4.54, 4.58] 4.37 [4.36, 4.38] 4.57 [4.56, 4.58]
III Experiments
This section describes the experimental setup of our two-example usage of the OpenACE benchmark. We focus on two particular use cases: objective evaluation of audio quality coding performance leveraging intrusive metrics and subjective evaluation through listening tests. In particular, we compare four DSP-based codecs: Opus [34], EVS [35], LC3 [36], and LC3+ [37]. Existing evaluation [38] reports LC3 performance on a test set of 24 German language samples only and misses assessment of the minimal supported bitrate of 16 kbps. Using the OpenACE, the size of the test increases by several orders of magnitude. The experiments only evaluated channel conditions without transmission errors, such as packet losses. The used signals are prepared by downmixing them to mono and converting the audio to a unified 16-bit PCM WAV format.
III-A Objective evaluation of audio quality vs bitrates
We leverage the intrusive VISQOL Audio [39] and POLQA [40] metrics for the objective evaluation. Given the 48 kHz sample rate requirement of the VISQOL algorithm, we select the full-band signals in the benchmark and resample them from 44.1 to 48 kHz using Kaiser window resampling. For objective metric computation, we truncate the audio to the first 10 seconds of each file. We evaluate the four coding algorithms at 16, 32, and 64 kbps.
Table III presents the objective scores for the various coding methods at the three different bitrates. The results show that at the low bitrate setting of 16 kbps, LC3+ shows the lowest performance, with LC3 performing slightly better. Notably, there is a significant gap of 0.4 between the two LC3-based codecs. EVS and Opus, on the other hand, achieve similar, higher results, with Opus having the highest score. Our results complement the previous LC3 assessment [38] that did not report the results obtained at 16 kbps. It is also fair to add that LC3 complexity encoding/decoding is significantly lower than Opus complexity. For example, in Hearing Aid music task, Opus has about twice as much complexity. This, however, results in degraded LC3 quality in lower bitrates.
For 32 kbps, we see an improvement in scores for Opus, LC3, and LC3+. Large quality increases were obtained for the LC3-based codecs, with LC3 and LC3+ gaining 0.49 and 0.89, respectively. On the other hand, Opus only obtained a slight increase of 0.07 but retained the best performance at 32 kbps. Surprisingly, the objective scores for EVS decreased slightly by increasing the bitrate from 16 to 32 kbps.
The results for the highest bitrate of 64 kbps show a significant increase in objective scores for all codecs. All methods obtain an approximate increase of 0.3 over the 32 kbps condition. At this bitrate, we can also observe that the two LC3-based codecs overtook EVS and Opus and obtained the highest performance of 4.39 MOS.
III-B Subjective evaluation of emotional speech at 16 kbps
The second experiment, conducted through subjective assessment, focused on studying the coding performance of emotional speech. To this end, we selected from the OpenACE a subset of five basic emotions: sadness, ecstasy, pain, fear, and anger. For reference, we added neutral (non-emotional) speech. The selected recordings were from the freeform section of the dataset. All of the tested codecs were set to operate at 16 kbps.
To evaluate the subjective quality of the different coding methods, we conduct an ITU-R BS.1534 MUSHRA test [41]. We used low- (3500 Hz) and mid-quality (7000 Hz) anchors generated using a type-1 Chebyshev filter. We truncated listening stimuli to the first 6 seconds.
TABLE IV: Mean MUSHRA scores of emotional speech encoding quality with 95% confidence intervals, obtained across all codecs.
Emotion Rating
Neutral 63.61 [61.26, 65.96]
Sadness 64.27 [61.95, 66.59]
Fear 63.14 [60.77, 65.51]
Pain 61.75 [59.32, 64.18]
Ecstasy 60.90 [58.50, 63.30]
Anger 60.51 [58.11, 62.91]
Fourteen assessors conducted the MUSHRA test through an online platform [^2] https://senselabonline.com/ using the headphones. Tab. IV shows obtained quality aggregated over all tested codecs. Type III ANOVA test revealed that there is a statistically significant difference in emotion encoding quality (F value = 7.4, $p<0.001$). We thus confirm that audio coding degrades signal quality, as reported by the EMO-Codec benchmark [42]. Their authors found that both DSP and ML-based codecs do not encode emotional speech sufficiently, which is more challenging in lower (below 24 kbps) bitrates. Besides, the used IEMOCAP dataset [43] needs to be licensed, which is not required with the OpenACE data.
Figure 1: Mean overall speech quality with 95% confidence intervals obtained of emotional speech encoding using the MUSHRA test [41].
Fig. 1 shows mean overall speech quality split across the encoding methods and emotional speech categories. The subjective test confirms the overall lower performances of LC3 and LC3+ codecs on the lower bitrate of 16 kbps. EVS usually provides better quality at the same bitrate as Opus [44], and we confirmed it also for emotional speech.
IV Discussion
The results for the objective assessment of the Opus, EVS, LC3, and LC3+ codecs on full-band audio highlight the differences between the codecs for varying bitrates. We observed that the LC3-based codecs performed significantly worse at lower bitrates than Opus and EVS. The results at 16 kbps also show that LC3+ performs worse than the standard LC3. This difference in measured performance is likely due to the difference in default behavior between the codecs, as the LC3+ codec applies a low-pass filter at 12 kHz, while LC3 applies a 20 kHz low-pass filter. Comparing the objective and subjective test results for LC3 and LC3+ reveals a discrepancy between the predicted objective scores and human evaluations. At 16 kbps, the objective results show a significant performance gap between LC3 and LC3+, which can not observed in the subjective test results, even when computing the VISQOL scores on only the subset used in the MUSRA test. This suggests that VISQOL may overemphasize the impact of lowpass filtering over human perception.
Audio coding has seen a paradigm shift to data-driven approaches, and neural encoding/decoding used in multimodal large language models will emphasize this research direction even more in the future. On the contrary, delay and complexity are intrinsic requirements for on-device processing, and further work will be done on ML-based codec models’ compression. All those developments can benefit from common benchmarking, such as the proposed OpenACE.
V Conclusions
We have described the audio and speech coding data that consist of 3 legacy coding data from the IEEE, ETSI and ITU-T, two open popular test data (VCTK and EBU SQAM), and two recent high-quality data released in 2024 (EARS and ODAQ). We have shown four DSP-based codecs benchmarking at 16 – 64 kbps and demonstrated their performance on emotional speech audio. The implemented open-source DSP-based codec evaluation can be easily used as the baseline method for current and future neural or hybrid codec R&D.
In the future, we want to maintain and extend the benchmark. Namely, we aim to add data augmentation like environmental noises and reverberation, and physical channel simulations for transmission error modeling, such as radio frequency interference. Such channel data augmentations may contribute to end-to-end quality/delay/complexity codec assessment. Adding multi-channel array(s) modeling will approach the codec benchmark to real-life conditions.